OpenStudy (anonymous):

Solve for x 4x-3+2x-9=2 ....I got confused because theres 2 x's i only know how to do it with one.

OpenStudy (anonymous):

Let's rewrite the equation like this: 4x + 2x - 3 - 9 = 2

OpenStudy (anonymous):

6... so 6x?

OpenStudy (anonymous):

now what do i do?

OpenStudy (anonymous):

4 + 2 = 6 4x + 2x = 6x same concept. So it would be 6x - 12 = 2 6x = 14 x = 7/3

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

happy to help

OpenStudy (anonymous):

hey, what do i do if the x's are on separate sides of the equal sign?

OpenStudy (texaschic101):

you either subtract one of them or add one of them...any way to get them both on the same side

OpenStudy (texaschic101):

example: 3x + 4 = 2x - 6 -- subtract 2x from both sides 3x - 2x + 4 = 2x - 2x - 6 -- combine like terms x + 4 = -6 -- subtract 4 from both sides x + 4 - 4 = -6 - 4 -- combine like terms x = -10 and you would just sub -10 in for x to solve for y

OpenStudy (anonymous):

so if the problem is Solve for X and the equation is 4x-3=x+18 what would i do?

OpenStudy (texaschic101):

oops...there is no y...lol..sorry about that

OpenStudy (texaschic101):

you would subtract x from both sides

OpenStudy (texaschic101):

let me show you.. 4x - 3 = x + 18 -- subtract x from both sides because what you do on one side of the equal sign, you have to do on the other 4x - x - 3 = x - x + 18 -- combine like terms (notice the x's cancel out o the right side) 3x - 3 = 18 -- now add 3 to both sides 3x - 3 + 3 = 18 + 3 -- combine like terms (the 3's cancel out on the right side) 3x = 20 x = 20/3 understand ?

OpenStudy (texaschic101):

oops..the 3's cancel out on the left side

OpenStudy (anonymous):

yes. thank you so much!

OpenStudy (texaschic101):

is it easier for you to see it this way.. 4x - 3 = x + 18 -x -x ===========add 3x - 3 = 18 +3 +3 ===========add 3x = 20 -- divide by 3 x = 20/3

OpenStudy (anonymous):

yea, it is. lol.

OpenStudy (texaschic101):

whatever is easiest for you to understand :)

OpenStudy (anonymous):

wait but wouldn't the 20 be 21 because its 18+3?

OpenStudy (texaschic101):

yep...your right...my bad

OpenStudy (texaschic101):

21/3 also reduces to 7

OpenStudy (anonymous):

its ok and one more question, sorry. at the very end is that the answer or would i still have to divide it by three to get the answer?

OpenStudy (texaschic101):

yes...if you have , for instance, 3x or 4x or any number times x, to get x by itself, you would have to divide

OpenStudy (anonymous):

k, thank you.

OpenStudy (texaschic101):

wait..you still have to divide by 3 giving you 21/3, which reduces to 7 you can even check your answer 4x - 3 = x + 18 4(7) - 3 = 7 + 18 28 - 3 = 25 25 = 25 (correct)

OpenStudy (anonymous):

all right, i got it. :)